OpenStudy (anonymous):

In 2001, Windsor, Ontario will receive its maximum amount of sunlight, 15.28 hrs, on June 21, and its least amount of sunlight, 9.08 hrs, on December 21. Due to the earth's revolution about the sun, the hours of daylight function is periodic. Determine an equation that can model the hours of daylight function for Windsor, Ontario. On what day(s) can Windsor expect 13.5 hours of sunlight?

OpenStudy (anonymous):

Dec-June= 12-6 =6 slope= (15.28-9.08)/ (12-6) = 1.03 y =1.03x+1, (1month ) 13.5 =1.03x+1 x = (13.5 -1)/1.03 x = 12.13 ≈ 12 days

OpenStudy (anonymous):

I have to go now